# Defending against Indirect Prompt Injection by Instruction Detection

### v1

@@ Abstract
The integration of Large Language Models (LLMs) with external sources is becoming increasingly common, with Retrieval-Augmented Generation (RAG) being a prominent example. However, this integration introduces vulnerabilities of Indirect Prompt Injection (IPI) attacks, where hidden instructions embedded in external data can manipulate LLMs into executing unintended or harmful actions. We recognize that the success of IPI attacks fundamentally relies in the presence of instructions embedded within external content, which can alter the behavioral state of LLMs. Can effectively detecting such state changes help us defend against IPI attacks? In this paper, we propose a novel approach that takes external data as input and leverages the behavioral state of LLMs during both forward and backward propagation to detect potential IPI attacks. Specifically, we demonstrate that the hidden states and gradients from intermediate layers provide highly discriminative features for instruction detection. By effectively combining these features, our approach achieves a detection accuracy of 99.60% in the in-domain setting and 96.90% in the out-of-domain setting, while reducing the attack success rate to just 0.12% on the BIPIA benchmark.
Figure 1: Overview of IPI attacks and our proposed method. (a) In a medical scenario of the RAG system, the LLM is misled by the external instruction embedded in a retrieved document to recommend company A’s medication. Our method performs instruction detection to defend against such attacks, removing such documents before they are passed to the LLM. (b) The success of IPI attacks lies in the presence of instructions embedded in external content, which alter the behavioral state of the LLM. Building on this insight, our method takes external data as input and pairs it with the response "Sure.", utilizing gradients and hidden states from optimally selected layers of the LLM as its behavioral state for instruction detection. (c) Our method achieves a detection accuracy of 99.60%, while reducing the ASR to just 0.12%.
Introduction
@@
Large language models (LLMs)  [1, 2, 3, 4, 5] have shown remarkable performance over various tasks, including question answering [6, 7], summarization [8, 9], and machine translation [10, 11]. Despite their impressive performance, LLMs often suffer from hallucinations [12, 13] and struggle with domain-specific or up-to-date knowledge, which limits their reliability in critical applications. To address these challenges, LLMs are increasingly integrated with external sources [14], a typical example being Retrieval-Augmented Generation (RAG) systems [15, 16]. This integration enables LLMs to generate responses that are more accurate, relevant, and temporally current, facilitating their applications in a wide range of domains.
@@
However, the inclusion of external content exposes LLMs to Indirect Prompt Injection (IPI) attacks. In such an attack, adversaries inject covert instructions into the external data retrieved by the system [17, 18, 19]. On the one hand, these hidden instructions may distort the retrieved information, leading the model to generate incorrect or misleading responses. On the other hand, they may cause the model to produce outputs that are entirely unrelated to the user’s intent, resulting in unexpected or irrelevant content. These vulnerabilities pose significant security and ethical risks, particularly in sensitive domains like healthcare [20, 21, 22], finance [23, 24], and legal systems [25, 26]. For example, as illustrated in Figure 1a, an instruction embedded in the external content could mislead the model into recommending medication from a specific company, even if it is not the most appropriate treatment for the patient, which results in harmful or biased medical advice.
@@
To mitigate the risk of IPI attacks, recent defenses have primarily focused on prevention [27, 28] by modifying prompts or fine-tuning models to ensure that LLMs adhere strictly to user instructions while ignoring external ones. However, detection [28], as an external method that enables proactively screening external resources to minimize time overhead and avoid the risk of affecting other benign inferences, remains underexplored and has yet to effectively detect IPI attacks. We recognize that the success of IPI attacks fundamentally lies in the presence of instructions embedded within external content, which alter the behavioral state of LLMs. Therefore, we hypothesize that this fundamental phenomenon—whether external data induces corresponding changes in the behavioral state of LLMs—can be leveraged to detect IPI attacks.
@@
Building on this insight, we first evaluate the effectiveness of hidden states and gradients from different layers of the LLM by employing them as features for instruction detection. Through experimentation on the validation set, we identify that the hidden states and gradients from intermediate layers consistently exhibit the best performance in differentiating normal external data from those containing hidden instructions. Specifically, we select the hidden states of the last token, as prior research indicates that the last token’s hidden state provides the most informative representation of the input sequence [29]. For the gradients, we focus on the gradients of self-attention layers, as previous studies suggest that self-attention layers capture the model’s behavioral characteristics, while feed-forward layers are more effective at encoding knowledge-based features [30, 31, 32]. Lastly, we fuse the hidden state features and the gradient features from the intermediate layer, which effectively integrates the complementary information captured by these two features. The fused features are then fed into a multi-layer perceptron (MLP) classifier, enabling effective detection of IPI attacks (Figure 1b).
@@
In our experiments, we consider normal external data as negative samples (without hidden instructions) and generate positive samples (with hidden instructions) by randomly inserting instructions into the negative samples. The external datasets include Wikipedia and News Articles, while the instruction data come from LaMini-instruction and BIPIA. Our approach achieves a detection accuracy of 99.60% in the in-domain setting and 96.90% in the out-of-domain setting, outperforming existing detection-based methods and several straightforward detection-based methods we propose. Furthermore, we conduct evaluation on the BIPIA benchmark (out-of-domain), where our method reduces the attack success rate (ASR) to just 0.12% (Figure 1c), surpassing the performance of the prevention-based methods reported in the benchmark. Overall, this work presents an effective detection-based defense against IPI attacks by leveraging the internal behavioral states of LLMs as discriminative signals, significantly improving the security of LLM-integrated systems. We hope our approach offers a practical solution to IPI attacks and provides novel insights into their underlying mechanisms and vulnerabilities.
Results
Datasets
@@ Results
In our experiments, we utilize external data from typical sources—Wikipedia [33] and News Articles [34]—while instructions come from LaMini-instruction [35] and BIPIA [27] datasets. Notably, there is no overlap between Wikipedia and News Articles, nor between LaMini-instruction and BIPIA, and they each belong to entirely different types and distributions of data.
Wikipedia
The dataset is constructed using Wikipedia dump files. Each data instance comprises the content of an entire Wikipedia article. In addition, we remove the overly long articles to ensure that they are not truncated during processing.
News Articles
The dataset contains 3,824 news articles, each featuring metadata including the title, subtitle, content, and publication date, sourced from multiple media outlets. Similarly, we remove the overly long articles to ensure that they are not truncated during processing.
LaMini-instruction
The dataset consists of 2.58 million pairs of instructions and corresponding responses, generated using GPT-3.5-Turbo, drawing from a wide range of existing resources of prompts, including Self-Instruct, P3, FLAN, and Alpaca.
BIPIA
BIPIA is the first benchmark aimed at evaluating the risk of IPI attacks on LLMs, and we use its instruction dataset for our experiments. The dataset consists of 15 attack types, categorized into task-irrelevant, task-relevant, and targeted attacks, with 5 instructions per attack type, resulting in a total of 75 instructions across both the training and test sets. These instructions were semi-automatically generated with the assistance of ChatGPT and manually reviewed for rationality.
Baselines
Our experiments involve two primary categories of baselines: detection-based defenses and prevention-based defenses. Detection-based defenses primarily focus on identifying IPI attacks, while prevention-based defenses, on the other hand, focus on ensuring that LLMs follow user instructions while ignoring external ones.
Detection-based Defenses
LLM (Zero-shot) [28]
Response Check [28]
TaskTracker [36]
LLM (Few-shot)
To enhance the performance of Naive LLM-based Detection, we attempt to leverage in-context learning [37] to strengthen the model’s capability to detect hidden instructions, where task demonstrations are integrated into the prompt in a natural language format.
LLM (Fine-tuning)
Prevention-based Defenses
In-context Learning [27]
Employ in-context learning to enable the model to distinguish between external data and user instructions, by providing samples where the model responds to input containing external data without being misled by the instruction embedded within the external data.
Multi-turn Dialogue [27]
Adversarial Training [27]
Incorporate adversarial learning during the LLM’s self-supervised fine-tuning phase, training the model to disregard instructions embedded within external content. The approach further adapts the model’s embedding layer to explicitly demarcate external content boundaries, enabling clearer distinction between external content and user instructions within inputs.
Experimental Setup
Our Method
@@
In our method, we utilize Llama-3.1-8B-Instruct [38] to extract behavioral states during its forward and backward propagation processes. Specifically, when extracting gradients as features, we pair the input external data with the response "Sure" as the typical reply to instructions. The extracted features are fed into an MLP classifier with hidden layer sizes set to (1024, 256, 64, 16). For training, we employ a dataset of 200 samples, evenly divided into 100 positive samples (with hidden instructions) and 100 negative samples (without hidden instructions). The balanced dataset ensures that the model learned to distinguish instructions effectively without being biased toward one class.
Detection Accuracy Comparison
@@
To compare our method with other detection-based defenses, we use a combination of external datasets and instruction datasets to create both positive and negative samples for instruction detection. Negative samples (without hidden instructions) are derived from external datasets, including Wikipedia and News Articles, as described earlier. Positive samples (with hidden instructions) are generated by modifying negative samples through the insertion of instructions sourced from LaMini-instruction and BIPIA datasets. The inserted instructions may appear at different positions in the text (e.g., at the beginning, middle, or end) to increase variability. We consider four distinct combinations of external datasets and instruction datasets: (1) Wikipedia with LaMini-instruction, (2) News Articles with LaMini-instruction, (3) Wikipedia with BIPIA, and (4) News Articles with BIPIA. For training and validation, we use the combination of Wikipedia and LaMini-instruction. For evaluation, we test our method on all four combinations of datasets, with each combination containing 2,000 samples. Among them, Wikipedia with LaMini-instruction is considered in-domain, while the other three combinations are out-of-domain to varying degrees. Notably, News Articles with BIPIA represent the highest level of out-of-domain shift. Therefore, when referring to out-of-domain performance in this paper, we specifically report results based on evaluations on News Articles with BIPIA. This diverse evaluation setup provides valuable insights into the reliability and generalizability of our method under different scenarios.
Attack Success Rate Comparison
To compare our method with other prevention-based defenses, we evaluate its impact on the ASR in the BIPIA [27] benchmark. We use GPT-3.5-Turbo [39] to assess whether the injected instructions within the external content lead the LLM to produce responses that deviate from the intended response, yielding the ASR. Specifically, we first apply our instruction detection method to the external data. Any external data for which no instructions are detected are subsequently used to conduct attacks. ASR is then computed by dividing the number of successful attack executions by the total sample count. To ensure comprehensive evaluation, we conduct IPI attack experiments on both an open-access model, Vicuna-7B [40], and a proprietary model, GPT-3.5-Turbo. This dual-model configuration enables us to assess the effectiveness of our approach across different architectures and access levels. Notably, our instruction detection method is trained on the combination of Wikipedia and LaMini-instruction, which have no overlap with the dataset used in the BIPIA benchmark. This evaluation thus offers an alternative perspective on the effectiveness of our method in defending against IPI attacks.
Overall Results
Detection Accuracy Comparison
@@
Table 1: Detection accuracy comparison of our proposed method and baseline approaches. The highest detection accuracy is indicated in bold. Here, ID denotes the in-domain setting, whereas OOD denotes the out-of-domain setting.
The effectiveness of our method is first evaluated through comparison with several detection-based defense methods, including existing approaches such as naive LLM (Zero-shot), Response Check, and TaskTracker, as well as several straightforward methods we propose to strengthen the model’s capability to detect hidden instructions: in-context learning and fine-tuning. As shown in Table 1, our method achieves superior performance over all baseline methods across all dataset combinations. LLM (Zero-shot), which directly queries the model, exhibits almost no capability to identify hidden instructions. Response Check, which evaluates the alignment of LLM outputs with intended tasks, provides moderate detection accuracy but is less effective overall, possibly because the inserted instructions do not necessarily alter the task corresponding to the response, making misalignment harder to detect. TaskTracker, which detects IPI attacks by contrasting the LLM’s activations before and after feeding the external data, achieves relatively high accuracy in the in-domain setting but remains less effective than our approach; additionally, its generalization capability is notably weaker. In-context learning, which provides task demonstrations within the prompt, offers minimal improvement over the naive approach, suggesting that simple prompting techniques are insufficient for enabling LLMs to detect hidden instructions. Fine-tuning LLMs significantly improves detection performance, but the method underperforms compared to our approach and exhibits weaker generalization across datasets, likely due to its inherent tendency to overfit specific training data rather than fully capturing the changes in the model’s behavioral state caused by hidden instructions. By leveraging discriminative features from intermediate layers, our method achieves superior performance and robust generalization, making it highly effective across diverse scenarios.
Attack Success Rate Comparison
GPT-3.5-Turbo Vicuna-7B
No Defense 33.57% 24.06%
In-context Learning 24.42% 16.85%
Multi-turn Dialogue 22.35% 14.66%
Adversarial Training - 0.52%
Ours 0.12% 0.03%
@@
Table 2: Comparison of ASR between our proposed method and baseline approaches. The lowest ASR is indicated in bold.
To assess the effectiveness of our approach in lowering ASR, we compare it with several prevention-based defenses, including in-context learning, multi-turn dialogue, and adversarial training. As illustrated in Table 2, our method consistently yields the lowest ASR on both open-access and proprietary models. Among the baseline methods, in-context learning and multi-turn dialogue, which are both black-box approaches, exhibit limited effectiveness in reducing ASR on both open-access and proprietary models, with ASR remaining significantly higher than that of our method. This indicates that simple structural modifications or prompting strategies fail to provide robust protection against IPI attacks. Adversarial training, a white-box method, demonstrates greater effectiveness in lowering ASR compared to black-box approaches. However, it still underperforms compared to our method and has limitations, especially for proprietary models, since it involves changes to the embedding layer and necessitates model fine-tuning. Our approach stands out for its ability to achieve superior ASR reduction while maintaining compatibility with both open-access and proprietary models, demonstrating its practicality and robustness against IPI attacks.
Ablation Study
Solely Utilizing Hidden States/Gradients

### v2

@@ Abstract
The integration of Large Language Models (LLMs) with external sources is becoming increasingly common, with Retrieval-Augmented Generation (RAG) being a prominent example. However, this integration introduces vulnerabilities of Indirect Prompt Injection (IPI) attacks, where hidden instructions embedded in external data can manipulate LLMs into executing unintended or harmful actions. We recognize that IPI attacks fundamentally rely on the presence of instructions embedded within external content, which can alter the behavioral states of LLMs. Can the effective detection of such state changes help us defend against IPI attacks? In this paper, we propose InstructDetector, a novel detection-based approach that leverages the behavioral states of LLMs to identify potential IPI attacks. Specifically, we demonstrate the hidden states and gradients from intermediate layers provide highly discriminative features for instruction detection. By effectively combining these features, InstructDetector achieves a detection accuracy of 99.60% in the in-domain setting and 96.90% in the out-of-domain setting, and reduces the attack success rate to just 0.03% on the BIPIA benchmark. The code is publicly available at https://github.com/MYVAE/Instruction-detection.
Defending against Indirect Prompt Injection by Instruction Detection
Tongyu Wen1thanks: Joint First Authors, Chenglong Wang211footnotemark: 1, Xiyuan Yang311footnotemark: 1, Haoyu Tang4, Yueqi Xie5, Lingjuan Lyu6, Zhicheng Dou1, Fangzhao Wu 7 1Renmin University of China, 2Peking University Shenzhen Graduate School, 3Wuhan University, 4University of Science and Technology of China, 5Hong Kong University of Science and Technology, 6Sony AI, 7Microsoft Research Asia Correspondence: yxieay@connect.ust.hk, dou@ruc.edu.cn, fangzwu@microsoft.com
1 Introduction
@@
Large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Brown et al., 2020; Chowdhery et al., 2023) have shown remarkable performance over various tasks, including question answering (Kamalloo et al., 2023; Singhal et al., 2023), summarization (Tang et al., 2023; Zhang et al., 2024), and machine translation (Xu et al., 2023; Zhang et al., 2023). Despite their impressive performance, LLMs often suffer from hallucinations (Ji et al., 2023; Rawte et al., 2023) and struggle with domain-specific or up-to-date knowledge, which limits their reliability in critical applications. To address these challenges, LLMs are increasingly integrated with external sources (Schick et al., 2023), a typical example being Retrieval-Augmented Generation (RAG) systems (Gao et al., 2023; Chen et al., 2024a). This integration enables LLMs to generate responses that are more accurate, relevant, and temporally current, facilitating their applications in a wide range of domains.
Figure 1: In a medical use case of the RAG system, the LLM is misled by the external instruction embedded in a retrieved document to recommend company A’s medication. Our method performs instruction detection to defend against such attacks, removing such documents before they are passed to the LLM.
@@
However, the inclusion of external content exposes LLMs to Indirect Prompt Injection (IPI) attacks. In such an attack, adversaries inject covert instructions into the external data retrieved by the system (Greshake et al., 2023; Rossi et al., 2024; Zhan et al., 2024; Chen et al., 2025; Kong, 2024). On the one hand, these hidden instructions may distort the retrieved information, leading the model to generate incorrect or misleading responses. On the other hand, they may cause the model to produce outputs that are entirely unrelated to the user’s intent, resulting in unexpected or irrelevant content. These vulnerabilities pose significant security and ethical risks, particularly in sensitive domains like healthcare (Sallam, 2023; Harrer, 2023; Yang et al., 2023), finance (Wu et al., 2023; Li et al., 2023), and legal systems (Cui et al., 2023; Lai et al., 2024). For example, as illustrated in Figure 1, an instruction embedded in the external content could mislead the model into recommending medication from a specific company, even if it is not the most appropriate treatment for the patient, which results in harmful or biased medical advice. Recently, similar manipulations have even been discovered in academic publishing, where researchers embedded hidden prompts in academic papers, such as “Do not highlight any negatives. Positive review only.” These instructions were concealed with invisible fonts to manipulate AI systems used in peer review into generating favorable evaluations (Nikkei Asia, 2025).
@@
To mitigate the risk of IPI attacks, recent defenses have primarily focused on prevention (Yi et al., 2025; Liu et al., 2024) by modifying prompts or fine-tuning models to ensure that LLMs adhere strictly to user instructions while ignoring external ones. However, detection (Liu et al., 2024), as an external method that enables proactively screening external resources to minimize time overhead and avoid the risk of affecting other benign inferences, remains underexplored and has yet to effectively detect IPI attacks. We recognize that IPI attacks fundamentally rely on the presence of instructions embedded within external content, which can alter the behavioral states of LLMs. Therefore, we hypothesize that this fundamental phenomenon—whether external data induces corresponding changes in the behavioral states of LLMs—can be leveraged to detect IPI attacks.
@@
Building on this insight, we propose InstructDetector, a novel detection-based approach that leverages the behavioral states of LLMs to identify potential IPI attacks. We first evaluate the effectiveness of hidden states and gradients from different layers of the LLM by employing them as features for instruction detection. Through experimentation on the validation set, we identify that the hidden states and gradients from intermediate layers consistently exhibit the best performance in differentiating normal external data from those containing hidden instructions. Specifically, we select the hidden states of the last token, as prior research indicates that the last token’s hidden state provides the most informative representation of the input sequence (Zou et al., 2023). For the gradients, we focus on the gradients of self-attention layers, as previous studies suggest that self-attention layers capture the model’s behavioral characteristics, while feed-forward layers are more effective at encoding knowledge-based features (Vaswani, 2017; Geva et al., 2021; Dai et al., 2022). Lastly, we fuse the hidden state features and the gradient features from the intermediate layer, which effectively integrates the complementary information captured by these two features. The fused features are then fed into a multi-layer perceptron (MLP) classifier, enabling effective detection of IPI attacks.
@@
In our experiments, we consider normal external data as negative samples (without hidden instructions) and generate positive samples (with hidden instructions) by randomly inserting instructions into the negative samples. The external datasets include Wikipedia and News Articles, while the instruction data come from LaMini-instruction and BIPIA. InstructDetector achieves a detection accuracy of 99.60% in the in-domain setting and 96.90% in the out-of-domain setting, outperforming existing detection-based methods and several straightforward detection-based methods we propose. Furthermore, we conduct evaluation on the BIPIA benchmark (out-of-domain), where our method reduces the attack success rate (ASR) to just 0.03%, surpassing the performance of the prevention-based methods reported in the benchmark. The contributions of our work can be outlined as follows:
  • We propose InstructDetector, a novel detection-based approach to defend IPI attacks, which leverages the internal behavioral states of LLMs as discriminative signals.
  • We find that hidden states and gradients from the intermediate layers of LLMs provide highly discriminative features for instruction detection.
  • Experiments demonstrate that InstructDetector achieves superior detection accuracy in both in-domain and out-of-domain settings, while significantly reducing the ASR compared to existing defense methods.
2 Related Work
2.1 Indirect Prompt Injection Defense
Defending against IPI attacks is a critical research area to ensure the secure and reliable use of LLMs (Greshake et al., 2023; Rossi et al., 2024; Zhan et al., 2024). Existing defenses are generally classified into prevention-based defences and detection-based defences (Yi et al., 2025; Liu et al., 2024).
Prevention-based defenses primarily focus on ensuring LLMs follow user instructions while ignoring external ones. These approaches are further divided into black-box defenses and white-box defenses. Black-box defenses (Yi et al., 2025; Hines et al., 2024; Wang et al., 2024; Wu et al., 2024a; Jia et al., 2024; Zhu et al., 2025) typically aim to isolate user instructions from external data, using carefully designed prompts to ensure LLMs disregard any hidden instructions within the external data. These methods work without access to the internal parameters of the model, focusing on input preprocessing and separation mechanisms. In contrast, white-box defenses (Yi et al., 2025; Chen et al., 2024b; Wang et al., 2025) utilize the internal parameters of the model and involve fine-tuning LLMs with samples of IPI attacks. By training on a diverse set of IPI scenarios, these methods enhance the robustness of LLMs to ignore external instructions while maintaining performance on the intended task.
Detection-based defenses, though relatively underexplored, aim to identify IPI attacks and can be generally divided into three main strategies. LLM (Zero-shot) (Liu et al., 2024; Chen et al., 2025) directly uses LLMs to identify hidden instructions in external data. Response Check (Liu et al., 2024) evaluates whether the model’s outputs remain consistent with the intended task. TaskTracker (Abdelnabi et al., 2024) detects IPI attacks by contrasting the LLM’s activations before and after feeding the external data, which indicates whether the user’s instruction is distorted by the instruction hidden in the external data. InstructDetector also falls under detection-based defenses, bridging the gap with a more robust mechanism for instruction detection.
2.2 Behavioral States of Large Language Models
Recent studies (Zou et al., 2023; Xie et al., 2024) have explored the internal mechanisms of LLMs, identifying hidden states and gradients as highly informative features for understanding and controlling their behavior. These behavioral states are increasingly recognized for their potential to enhance the transparency and safety of LLMs.
Hidden states, especially those from intermediate layers, have been shown to encode rich and insightful representations of given inputs. RepE (Zou et al., 2023) utilizes representations from the last token’s hidden states to monitor and manipulate high-level cognitive phenomena in LLMs. Furthermore, a recent study (Skean et al., 2024) has explored the effectiveness of intermediate features across different LLM architectures, revealing that intermediate features often yield richer information than final-layer for downstream use.
Gradients provide another critical lens for analyzing the LLM’s behavior. Gradsafe (Xie et al., 2024) leverages the observation that adversarial prompts generate distinct gradient patterns compared to safe prompts, enabling effective jailbreak prompts detection without additional training by analyzing gradients related to safety-critical parameters. Additionally, much literature (Vaswani, 2017; Geva et al., 2021; Dai et al., 2022) has explored the functions of self-attention layers and feed-forward layers, providing insights into where to focus when analyzing gradients in our work. Research has shown that self-attention layers capture behavioral characteristics, such as linguistic dependencies and token relationships, while feed-forward layers encode knowledge-based features, enabling the model to leverage the knowledge learned during training. As the instruction recognition task primarily relies on the behavioral characteristics of LLMs, we focus on the gradients of self-attention layers.
3 Methodology
Figure 2: IPI attacks fundamentally rely on the presence of instructions embedded in external content, which can alter the behavioral states of LLMs. Building on this insight, InstructDetector takes external data as input and pairs it with the response "Sure.", utilizing gradients and hidden states from optimally selected layers of the LLM as its behavioral states for instruction detection.
3.1 Overview
In our proposed method, InstructDetector, we aim to detect IPI attacks through the behavioral states of LLMs, hypothesizing that changes in the behavioral states of LLMs induced by embedded instructions in external content can be effectively utilized to detect such attacks. To achieve this, we fuse the hidden states and gradients from the most effective layers, integrating complementary information captured by both features. These fused features are then fed into an MLP classifier, enabling accurate and robust detection of IPI attacks. The overall framework is illustrated in Figure 2, with detailed processes discussed in the following sections.
3.2 Hidden States Extraction
To leverage hidden states as features, we first take external data as the input of the LLM and extract the hidden states corresponding to the last token at each layer. These hidden states are then fed into an MLP classifier to assess their ability to distinguish between normal external data and those containing hidden instructions. InstructDetector uses the Llama-3.1-8B-Instruct model, which consists of 32 layers. Through experimentation on the validation set, we identify that the hidden states from the 14th layer provide the best performance in instruction detection. Therefore, we select the last token’s hidden state from the 14th layer, a vector with a dimension of 4096, as the first input of the feature fusion module.
3.3 Gradients Extraction
To leverage gradients as features, we first take external data as the input of the LLM, paired with a typical response to instructions, such as "Sure," and compute the gradients for the model parameters at each layer during back propagation. Based on prior research indicating that self-attention layers capture the model’s behavioral characteristics, while the feed-forward layers are more effective at encoding knowledge-based features, we concentrate on the gradients of self-attention layers. Experimental results on the validation set demonstrate that the gradients from the 14th layer, consistent with the layer identified for hidden states, yield the best performance in distinguishing between normal external data and those with hidden instructions.
3.4 Feature Fusion
4 Experiment
4.1 Datasets
@@
In our experiments, we utilize external data from typical sources—Wikipedia (Foundation, ) and News Articles (dai, 2017)—while instructions come from LaMini-instruction (Wu et al., 2024b) and BIPIA (Yi et al., 2025) datasets. Notably, there is no overlap between Wikipedia and News Articles, nor between LaMini-instruction and BIPIA, and they each belong to entirely different types and distributions of data. Detailed descriptions of each dataset are provided in Appendix A.1.
4.2 Baselines
Our experiments involve two primary categories of baselines: detection-based and prevention-based defenses. Detection-based defenses primarily focus on identifying IPI attacks. These include LLM (Zero-shot) (Liu et al., 2024) and LLM (Few-shot), which directly query the LLM to identify if there is any hidden instruction within the external content in a zero-shot or few-shot setting; Response Check (Liu et al., 2024), which checks whether the response aligns with the intended task; TaskTracker (Abdelnabi et al., 2024), which contrasts the LLM’s activations before and after feeding the external data; and LLM (Fine-tuning), which conducts supervised fine-tuning using task-specific annotated data.
Prevention-based defenses, on the other hand, focus on ensuring that LLMs follow user instructions while ignoring external ones. Strategies include Multi-turn Dialogue (Yi et al., 2025), which separates user prompts from external data using multi-turn dialogue; In-context Learning (Yi et al., 2025), which employs in-context learning to teach the model how to resist misleading input patterns; and Adversarial Training (Yi et al., 2025), which applies adversarial training to help the model distinguish and ignore instruction-carrying content from external sources.
A detailed description of each baseline method can be found in Appendix A.2, and the implementation details and configurations are provided in Appendix A.3.
4.3 Experimental Setup
4.3.1 InstructDetector
@@
InstructDetector utilizes Llama-3.1-8B-Instruct (Dubey et al., 2024) to extract behavioral states during its forward and backward propagation processes. Specifically, when extracting gradients as features, we pair the input external data with the response "Sure" as the typical reply to instructions. The extracted features are fed into an MLP classifier with hidden layer sizes set to (1024, 256, 64, 16). For training, we employ a dataset of 200 samples, evenly divided into 100 positive samples (with hidden instructions) and 100 negative samples (without hidden instructions). The balanced dataset ensures that the model learned to distinguish instructions effectively without being biased toward one class.
4.3.2 Detection Accuracy Comparison
@@
To compare InstructDetector with other detection-based defenses, we use a combination of external datasets and instruction datasets to create positive and negative samples for instruction detection. Negative samples are derived from external datasets, and positive samples are generated by randomly inserting instructions into negative samples. For training and validation, we use the combination of Wikipedia and LaMini-instruction. For evaluation, we test methods on all four combinations of datasets, with each combination containing 2,000 samples. Among them, Wikipedia with LaMini-instruction is considered in-domain, while the other three combinations are out-of-domain to varying degrees. Notably, News Articles with BIPIA represent the highest level of out-of-domain shift. Therefore, when referring to out-of-domain performance in this paper, we specifically report results based on evaluations on News Articles with BIPIA.
@@
Table 1: Detection accuracy comparison of InstructDetector and baseline approaches. The highest detection accuracy is indicated in bold. Here, ID denotes the in-domain setting, whereas OOD denotes the out-of-domain setting.
GPT-3.5-Turbo GPT-4o Vicuna-7B Qwen2.5-7B-Instruct
No Defense 33.57% 39.68% 24.06% 35.60%
In-context Learning 24.42% 33.00% 16.85% 30.73%
Multi-turn Dialogue 22.35% 14.08% 14.66% 16.19%
Adversarial Training - - 0.52% 0.68%
InstructDetector 0.12% 0.13% 0.03% 0.10%
@@
Table 2: Comparison of ASR between InstructDetector and baseline approaches. The lowest ASR is indicated in bold.
4.3.3 Attack Success Rate Comparison
To compare InstructDetector with other prevention-based defenses, we evaluate its impact on the ASR in the BIPIA (Yi et al., 2025) benchmark. We use GPT-3.5-Turbo (Dale, 2021) to assess whether the injected instructions within the external content lead the LLM to produce responses that deviate from the intended response, yielding the ASR. Specifically, we first apply our instruction detection method to the external data. Any external data for which no instructions are detected are subsequently used to conduct attacks. ASR is then computed by dividing the number of successful attack executions by the total sample count. To ensure comprehensive evaluation, we conduct IPI attack experiments on both an open-access model, Vicuna-7B (Chiang et al., 2023) and Qwen2.5-7B-Instruct (Bai et al., 2023), as well as proprietary models, GPT-3.5-Turbo and GPT-4o. Notably, our instruction detection method is trained on the combination of Wikipedia and LaMini-instruction, which have no overlap with the dataset used in the BIPIA benchmark.
4.4 Overall Results
4.4.1 Detection Accuracy Comparison
The effectiveness of InstructDetector is first evaluated through comparison with several detection-based defenses, including existing approaches such as naive LLM (Zero-shot), Response Check, and TaskTracker, as well as several straightforward methods we propose to strengthen the model’s capability to detect hidden instructions: in-context learning and fine-tuning. As shown in Table 1, InstructDetector achieves superior performance over all baselines across all dataset combinations.
LLM (Zero-shot), which directly queries the model, exhibits almost no capability to identify hidden instructions. Response Check, which evaluates the alignment of LLM outputs with intended tasks, provides moderate detection accuracy but is less effective overall, possibly because the inserted instructions do not necessarily alter the task corresponding to the response, making misalignment harder to detect. TaskTracker, which detects IPI attacks by contrasting the LLM’s activations before and after feeding the external data, achieves relatively high accuracy in the in-domain setting but remains less effective than InstructDetector; also, its generalization capability is notably weaker. A further comparative analysis between InstructDetector and TaskTracker is provided in Appendix A.3.
In-context learning, which provides task demonstrations within the prompt, offers minimal improvement over the naive approach, suggesting that simple prompting techniques are insufficient for enabling LLMs to detect hidden instructions. Fine-tuning LLMs significantly improves detection performance, but the method underperforms compared to InstructDetector and exhibits weaker generalization across datasets, likely due to its inherent tendency to overfit specific training data rather than fully capturing the changes in the model’s behavioral states caused by hidden instructions.
By leveraging discriminative features from intermediate layers, InstructDetector achieves superior performance and robust generalization, making it highly effective across diverse scenarios.
4.4.2 Attack Success Rate Comparison